# Are You There God? Lightweight Narrative Annotation of Christian Fiction with LMs

### v1

@@ 4.3 Characterizing Acts of God
To further characterize the acts of God identified by the classification system described above, we design two further prompts for Gemma 3n. These prompts seek to identify who is affected by each act of God (individuals or groups) and whether God’s action is loving, punishing, both, or neutral. Again, we begin with prompts written for human annotators and transform them for use with Gemma 3n using GPT-4o. Both prompts take as input the descriptions of God’s acts output by the primary classifier. Final versions of both prompts can be seen in Appendix B.2.
@@ 5 Characterizing Christian Fiction
We focus on two primary trends in our topic modeling results: expression of faith and portrayal of gender. [^10] A full list of topics and associated keywords can be found in Appendix C.1. We identify four faith-related topics: Confessional Prayer (#13: believe, father, please, forgive), Glorifying God (#43: God, heaven, salvation, amen), Apocalyptic Faith (#48: Israel, tribulation, death, messiah), and Congregation Worship (#63: church, Sunday, music, sermon). Surprisingly, these topics are comparatively rare despite the overall importance of faith in the genre (Appendix C.2). Glorifying God and Apocalyptic Worship are highly positively correlated (Pearson’s R: 0.84, $p<10^{-24}$) whereas Confessional Prayer and Congregation worship are slightly negatively correlated (Pearson’s R: -0.26, $p<0.05$); these results suggest that three distinct faith styles are employed by authors.
We also identify five topics tied to the “Cult of Domesticity,” the Evangelical Christian belief that a woman’s God-given role is as a mother and homemaker: Food and Cooking (#1: food, kitchen, meal, hungry), Family Members (#4: family, father, children, parents), Family Relationships (#8: mom, dad, father, kids), Motherhood and Pregnancy (#27: baby, mother, pregnant, birth) and Cleaning the Home (#34: clean, bathroom, towel, wet). All share positive significant correlations with at least two other topics in the grouping; Food and Cooking is positively correlated with every other domestic topic. The close relationship between these topics suggests that Christian Fiction depicts an idealized lifestyle for women that emphasizes domesticity [13].
@@
These topic groupings provide evidence that the Left Behind series differs considerably from Christian Fiction more broadly. Although there is no significant difference between the prevalence of the Confessional Prayer and Congregation Worship topics in the Left Behind books and the broader corpus, both the Glorifying God (4.51 vs. 1.38, $p<10^{-17}$) and Apocalyptic Faith (5.39 vs. 0.41, $p<10^{-37}$) topics were significantly more dominant in the Left Behind books. [^11] Independent t-tests for difference in means were performed using sklearn. The Left Behind series also featured domestic themes at significantly lower frequencies than the rest of the corpus.121212#1: 0.45 vs. 1.75, ($p<10^{-5}$), #4: 1.16 vs. 2.02 ($p<10^{-3}$), #8: 0.57 vs. 1.32 ($p<0.05$), #27: 0.41 vs. 0.65 ($p<0.05$), #34: 0.37 vs. 0.77 ($p<10^{-4}$) It is unclear whether these differences stem from Left Behind’s apocalyptic setting or gendered preferences in writing.
To pursue this question, we examine whether these topics appear to significantly different extents in books written by female and male authors. To do this, we exclude all Left Behind novels and books co-written by a male and female author from the dataset. We find no significant differences between topic prominence for any of the four faith-focused topics. However, the Food and Cooking (1.99 vs. 1.08, $p<10^{-3}$) and Family Members (2.18 vs. 1.44, $p<10^{-3}$) topics were significantly more prominent on average in books written by women.
We identify three primary patterns in the topic model. First, the Left Behind series approaches faith in a distinct way; Apocalyptic Faith and Glorifying God are highly prevalent in these books while largely absent from the rest of the corpus. Second, we see that this difference in religious portrayal is not due to gender; male and female authors portray faith to similar extents. Third, we find that an author’s gender does shift the prevalence of gendered topics: women engage more with topics within the Cult of Domesticity. This makes clear the divide between the intensely religious themes of the Left Behind series and the more contemporary emphasis on domestic themes authored by women.
6 Identifying Acts of God
Of all the passages in our corpus, 24.9% or 5,086 are labeled as containing “acts of God.” On average, 26.1% of passages from each novel contain an act of God; however, this value ranges from 2.1% [^13] Auralia’s Colors by Jeffrey Overstreet to 82.6%  [^14] Glorious Appearing by Jerry B. Jenkings and Tim LaHaye across all novels (Figure 2). This demonstrates that, although these books are positioned in an explicitly religious subgenre often assumed to contain religious indoctrination, there is actually wide variation in how much authors portray acts of God.
We find that three of the four faith topics — Confessional Prayer (#13 – Pearson’s R: 0.28, $p<10^{-3}$), Glorifying God (#43 – Pearson’s R: 0.88, $p<10^{-30}$), and Apocalyptic Faith (#48 – Pearson’s R: 0.71, $p<10^{14}$) — significantly correlate with the frequency of divine action. These correlations hold even when the Left Behind novels are removed from the dataset.151515#13 – 0.38 ($p<10^{-4}$), #43 – 0.82 ($p<10^{-19}$), #48 – 0.55 ($p<10^{-7}$) However, no such significant relationship exists for the fourth faith topic, Congregation Worship, which is the group-oriented. This highlights that explicit actions of God occur more frequently in books which include more individualized expressions of faith, which aligns with American Evangelicalism’s hyper-individualistic tendencies.
@@ 6 Identifying Acts of God
We also see that acts of God are more densely clustered later in novels (Figure 3). On average, acts of God appear 56.2% of the way through a plot. There are several possible explanations for this. First, acts of God may simply be examples of important narrative events, which frequently occur towards the end of novels. Second, Christian Fiction authors may wish to defer “religious talk” to a point where readers are invested in the characters or plot, so as to avoid putting off secular readers. This approach serves to increase readership, but also serves a spiritual goal of introducing non-Christian readers to Christian values and viewpoints.
Finally, we see that there is no significant difference between the frequency with which male and female authors depict acts of God. However, a significantly higher proportion of passages from the Left Behind novels depict acts of God on average (46.6% vs. 22.85%, $p<10^{-9}$); [^16] All t-tests for difference in means are again performed with sklearn. this is clearly visible in Figure 2. Left Behind explicitly sets out to narrate events from the Book of Revelation and in doing so frequently portrays God as a core actor. In contrast, most books invoke God’s intervention more sparingly. Notably, the novel featuring acts of God with the second-highest frequency (Unashamed by Francine Rivers) is not from the Left Behind series but is a retelling of the story of Rahab from the Bible. We would expect books explicitly retelling stories from the Bible to feature more acts of God; their high ranking by our classifier therefore reinforces our confidence in its accuracy.
@@ 7 Characterizing Acts of God
Of all the acts of God identified in our corpus, 86.8% are labeled as affecting individuals and 13.2% affect groups. 79.4% of divine acts are classified as loving, 16.1% as both loving and punishing, 3.6% as punishing, and 0.9% as neutral. We find that the overwhelming majority of divine acts in our corpus are loving acts that affect individual people; this trend is notable even taking into account that our classifiers tend to over-predict these labels. Although the majority of acts directed at both individuals and groups are loving, a far greater proportion of acts impacting groups are punishing in nature (6.25% vs. 3.22%) or both loving and punishing (29.17% vs. 14.16%). This reveals a difference in how God’s actions are characterized depending on who is affected.
@@
The frequency of divine acts impacting individuals is negatively correlated with Apocalyptic Faith (#48 – Pearson’s R: -0.65, $p<10^{-12}$), which is intuitive given apocalyptic acts often affect groups. A negative correlation also exists with Glorifying God (#43 – Pearson’s R: -0.58, $p<10^{-9}$), which is surprising since this topic closely aligns with conventional Evangelical individualistic expressions of faith. There is also some connection between the characterization of divine acts and faith. The frequency of loving acts is negatively correlated with Apocalyptic Faith (#48 – Pearson’s R: -0.67, $p<10^{-13}$) and the frequency of acts that are both loving and punishing is positively correlated with Apocalyptic Faith (Pearson’s R: 0.72, $p<10^{-15}$) and Glorifying God (#43 – Pearson’s R: 0.65, $p<10^{-11}$). Combined, these trends demonstrate that apocalyptic themes and the glorification of God are associated with a greater frequency of punishing divine acts and acts of God which impact groups. [^17] Correlations hold without Left Behind books: #48 + Individuals – -0.62 ($p<10^{-9}$), #43 + Individuals – -0.23 ($p<0.05$), #48 + Loving – -0.32 ($p<10^{-3}$), #48 + Both – 0.37 ($p<10^{-4}$), #43 + Both – 0.27 ($p<0.05$)
@@
We again find that the Left Behind novels are unique within our corpus. On average, a significantly smaller proportion of divine acts in these books are targeted towards individuals (76.03% vs. 90.11%, $p<10^{-8}$, Figure 4), likely because they depict how God transforms the world in preparation for His Second Coming. We also see that a significantly smaller proportion of God’s actions in the Left Behind books are depicted as loving (63.24% vs. 84.94%, $p<10^{-4}$, Figure 5), while a significantly larger proportion of acts are both loving and punishing (28.73% vs. 11.31%, $p<10^{-11}$). Moreover, nearly twice the proportion of actions are punishing on average (7.00% vs. 2.81%, $p<10^{-3}$). These differences reinforce that the apocalyptic narrative of the Left Behind series is not typical of Christian Fiction as a genre and the books should not be used as a proxy for the genre as a whole.
Finally, we find no significant relationship between author gender and the depiction of divine acts affecting individuals even when the Left Behind novels are excluded from the dataset. Thus, we see no evidence that male and female authors differ in the frequency with which they portray acts of God or that they depict the scope of God’s actions differently. However, we do observe that a higher proportion of divine acts written by female authors are loving on average (86.10% vs. 80.01%, $p<0.05$) and a lower proportion are punishing (2.09% vs. 5.14%, $p<0.05$). This may be because female authors write more romance novels (Appendix A.3) which prominently feature God’s love as an essential part of the narrative [36]. It may also indicate how female authors center love and forgiveness as core elements of their spiritual identity [17].
@@ 8 Conclusion
Our multidisciplinary work suggests a combination of literary, cultural, and methodological insights. Scholarship has paid little attention to Christian Fiction published after the conclusion of the Left Behind series, instead focusing on the male-dominated cultural movements seeking to assert Evangelical Christianity through political power. However, the Christian Fiction community represents a fundamentally different approach: one where Christian women, organized around faith and fiction, seek to transform the individual hearts and minds of Americans. By examining God’s depiction in Christian Fiction, we gain access to a perspective on Evangelical faith that stands apart from its louder mainstream projection. Our results also confirm that the Left Behind series is fundamentally different from other Christian Fiction novels; it focuses more explicitly on faith, depicts acts of God far more frequently, and is more likely to characterize God as punishing. In addition, we demonstrate that within Christian Fiction female authors are more likely to feature topics related to the Cult of Domesticity and present a loving God.
From a methodological perspective, while it may be tempting — or terrifying — to think of new computational technologies as replacing scholarship with mere button pushing, we find that effectively using these methods requires an extensive, iterative process. More established methods such as topic modeling produce results of a clearly limited and general character that both facilitate and require interpretation. Newer prompt-based methods that both interpret text and produce human-readable outputs nevertheless require serious work before we understand what we want to ask of them and how we want to ask for it. However, combining the expertise of subject-matter experts and traditional qualitative research methods with powerful contemporary LMs allows for the analysis of complex, subtle concepts on previously unreachable scales, opening possibilities for many new avenues of research.
@@ Acknowledgements
We would like to thank Erin Boell, Carolyn Chui, Estefania DeJesus, Mia Ferrante, Alexandra Gier, Georgia Lawrence, Sushmi Majumder, and Elina Natarajan for their substantial annotation contributions. In addition, we thank Axel Bax, Federica Bologna, Anna Choi, Sil Hamilton, Kiara Liu, Landon Schnabel, Rosamond Thalken, Andrea Wang, Matthew Wilkens, and Shengqi Zhu for their thoughtful feedback. This work was supported in part by the NEH project AI for Humanists.

### v2

@@ 4.3 Characterizing Acts of God
To further characterize the acts of God identified by the classification system described above, we designed two further prompts for Gemma 3n. These prompts seek to identify who is affected by each act of God (individuals or groups) and whether God’s action is loving, punishing, both, or neutral. Again, we begin with prompts written for human annotators and transform them for use with Gemma 3n using GPT-4o. Both prompts take as input the descriptions of God’s acts output by the primary classifier. Final versions of both prompts can be seen in Appendix B.2.
@@ 5 Characterizing Christian Fiction
We focus on two primary trends in our topic modeling results: expression of faith and portrayal of gender. [^9] A full list of topics and associated keywords can be found in Appendix C.1. We identify four faith-related topics: Confessional Prayer (#13: believe, father, please, forgive), Glorifying God (#43: God, heaven, salvation, amen), Apocalyptic Faith (#48: Israel, tribulation, death, messiah), and Congregation Worship (#63: church, Sunday, music, sermon). Surprisingly, these topics are comparatively rare despite the overall importance of faith in the genre (Appendix C.2). Glorifying God and Apocalyptic Faith are highly positively correlated (Pearson’s R: 0.84, $p<10^{-21}$), [^10] All $p$-values reported have had Bonferroni correction applied using the statsmodels package to account for multiple similar statistical tests being run on the same dataset. but no significant correlations exist between the other faith categories; this suggests that distinct faith styles are employed by authors.
We also identify five topics tied to complementarianism, the theological argument that men and women were created to serve distinct roles, with men taking on leadership positions in the household and wider society while women primarily serve the family unit. These topics are Food and Cooking (#1: food, kitchen, meal, hungry), Family Members (#4: family, father, children, parents), Family Relationships (#8: mom, dad, father, kids), Motherhood and Pregnancy (#27: baby, mother, pregnant, birth), and Cleaning the Home (#34: clean, bathroom, towel, wet). Prevalence of the Cooking, Family Relationships, and Cleaning the Home topics are all significantly positively correlated as are Family Members and Motherhood and Pregnancy. The relationships between these topics suggests that Christian Fiction depicts an idealized lifestyle for Evangelical women that emphasizes domesticity [15].
These topic groupings provide evidence that the Left Behind series differs considerably from Christian Fiction more broadly. Although there is no significant difference between the prevalence of the Confessional Prayer and Congregation Worship topics in the Left Behind books and the broader corpus, both the Glorifying God (4.51 vs. 1.38, $p<10^{-17}$) and Apocalyptic Faith (5.39 vs. 0.41, $p<10^{-37}$) topics were significantly more dominant in the Left Behind books. [^11] Independent t-tests for difference in means were performed using sklearn The Left Behind series also features Cooking and Cleaning the Home at significantly lower frequencies than the rest of the corpus.121212#1: 0.45 vs. 1.75, ($p<0.01$), #34: 0.37 vs. 0.77 ($p<0.05$)) It is unclear whether these differences stem from Left Behind’s apocalyptic setting or gendered preferences in writing.
To pursue this question, we examine whether the prevalence of these topics differs significantly in books written by female and male authors. To do this, we exclude all Left Behind novels and books co-written by a male and female author from the dataset. We find no significant differences between topic prominence between male and female authors for any of the four faith-focused topics or any of the five topics tied to complementarianism. However, we do see that women authors depict Joyful Affection (#2: love, smile, kiss, joy) more on average (1.02 vs. 1.99, $p<0.05$) and Violence (#23: blood, body, gun, dead) less on average (5.30 vs. 3.16, $p<0.05$) than male authors, which aligns with standards of Evangelical femininity.
Overall, we identify three primary patterns in the topic model: the Left Behind series approaches faith in a distinct way, this difference in religious portrayal is not due to gender — male and female authors portray faith to similar extents, but an author’s gender does shift the prevalence of some gendered topics. This makes clear the divide between the intensely religious themes of the Left Behind series and more contemporary Christian Fiction.
6 Identifying Acts of God
@@
Of all the passages in our corpus, 24.9% or 5,086 are labeled as containing acts of God. On average, 26.1% of passages from each novel contain an act of God; however, this value ranges from 2.1% [^13] Auralia’s Colors by Jeffrey Overstreet to 82.6% [^14] Glorious Appearing by Jerry B. Jenkings and Tim LaHaye across all novels (Figure 2). This demonstrates that, although these books are positioned in an explicitly religious subgenre often assumed to contain religious indoctrination, there is actually wide variation in how frequently authors portray acts of God.
We find that two of the four faith topics — Glorifying God (Pearson’s R: 0.88, $p<10^{-28}$), and Apocalyptic Faith (Pearson’s R: 0.71, $p<10^{13}$) — significantly correlate with the frequency of divine action. When the Left Behind novels are removed from the dataset, these correlations hold151515#43 – 0.82 ($p<10^{-18}$), #48 – 0.55 ($p<10^{-5}$) and a significant relationship between the frequency of divine action and Confessional Prayer emerges.1616160.38, $p<0.05$ However, no such significant relationship exists for Congregation Worship. This highlights that explicit actions of God occur more frequently in books which emphasize individualized expressions of faith, which aligns with American Evangelicalism’s hyper-individualistic tendencies.
@@
We also see that acts of God are more densely clustered later in novels (Figure 3). On average, acts of God appear 56.2% of the way through a plot. There are two possible explanations for this. First, acts of God may simply serve as important narrative events, which frequently occur towards the end of novels. Second, Christian Fiction authors may wish to defer “religious talk” to a point where readers are invested in the characters or plot, so as to avoid putting off secular readers. This approach serves to increase readership, but also serves a spiritual goal of introducing non-Christian readers to Christian values and viewpoints.
Finally, we see that there is no significant difference between the frequency with which male and female authors depict acts of God. However, a significantly higher proportion of passages from the Left Behind novels depict acts of God on average (46.6% vs. 22.85%, $p<10^{-9}$); [^17] All t-tests for difference in means are again performed with sklearn with Bonferroni correction with statsmodels. this is clearly visible in Figure 2. Left Behind explicitly sets out to narrate events from the Book of Revelation and in doing so frequently portrays God as a core actor. In contrast, most books invoke God’s intervention more sparingly. Notably, the novel featuring acts of God with the second-highest frequency (Unashamed by Francine Rivers) is not from the Left Behind series but is a retelling of the story of Rahab from the Bible. We would expect books explicitly retelling stories from the Bible to feature more acts of God; their high ranking by our classifier therefore reinforces our confidence in its accuracy.
@@ 7 Characterizing Acts of God
We find that the overwhelming majority of divine acts in our corpus are loving acts that affect individuals; this trend is notable even taking into account that our classifiers tend to over-predict these labels. Of all the acts of God identified in our corpus, 86.8% are labeled as affecting individuals and 13.2% affect groups. 79.4% of divine acts are classified as loving, 16.1% as both loving and punishing, 3.6% as punishing, and 0.9% as neutral. Although the majority of acts directed at both individuals and groups are loving, a far greater proportion of acts impacting groups are punishing in nature (6.25% vs. 3.22%) or both loving and punishing (29.17% vs. 14.16%). This reveals a difference in how God’s actions are characterized depending on who is affected.
@@
The frequency of divine acts impacting individuals is negatively correlated with Apocalyptic Faith (Pearson’s R: -0.65, $p<10^{-10}$), which is intuitive given apocalyptic acts often affect groups. A negative correlation also exists with Glorifying God (Pearson’s R: -0.58, $p<10^{-7}$), which is surprising since this topic closely aligns with conventional Evangelical individualistic expressions of faith. There is also some connection between the characterization of divine acts and faith. The frequency of loving acts is negatively correlated with Apocalyptic Faith (Pearson’s R: -0.67, $p<10^{-11}$) and Glorifying God (Pearson’s R: -0.58, $p<10^{-7}$) and the frequency of acts that are both loving and punishing is positively correlated with both these topics (0.72, $p<10^{-13}$ — 0.65, $p<10^{-10}$). Combined, these trends demonstrate that apocalyptic themes and the glorification of God are associated with a greater frequency of punishing divine acts and acts of God which impact groups. [^18] Correlations largely do not hold without Left Behind books, largely due to the small number of group and punishing acts that appear outside of them.
@@
We again find that the Left Behind novels are unique within our corpus. On average, a significantly smaller proportion of divine acts in these books are targeted towards individuals (76.03% vs. 90.11%, $p<10^{-7}$, Figure 4), likely because they depict how God transforms the world in preparation for His Second Coming. We also see that a significantly smaller proportion of God’s actions in the Left Behind books are depicted as loving (63.24% vs. 84.94%, $p<10^{-10}$, Figure 5), while a significantly larger proportion of acts are both loving and punishing (28.73% vs. 11.31%, $p<10^{-11}$). Moreover, nearly twice the proportion of actions are punishing on average (7.00% vs. 2.81%, $p<0.05$). These differences reinforce that the apocalyptic narrative of the Left Behind series is not typical of Christian Fiction as a genre and the books should not be used as a proxy for the genre as a whole. Finally, we find no significant evidence that male and female authors differ in the frequency with which they portray acts of God or how they characterize these acts.
@@ 8 Conclusion
Our multidisciplinary work provides a combination of literary, cultural, and methodological insights. Scholarship has paid little attention to Christian Fiction published after the conclusion of the Left Behind series, instead focusing on the male-dominated cultural movements seeking to assert Evangelical Christianity through cultural power. In contrast, the Christian Fiction community has adopted a fundamentally different approach: one where Christian women, organized around faith and fiction, seek to transform the individual hearts and minds of Americans. This work begins the project of describing God’s portrayal in Christian Fiction, and thereby adds to our understanding of Evangelical ministry and faith. Ultimately, we find that the Left Behind series fundamentally differs from predominant Christian Fiction novels due to its depiction of a more active and punishing God while the authors who dominate Christian Fiction circles today largely write stories that model Evangelical femininity and depict a loving God.
From a methodological perspective, while it may be tempting — or terrifying — to think of new computational technologies as replacing scholarship with mere button pushing, we find that effectively using these methods requires an extensive, iterative process. More established methods such as topic modeling produce results of a clearly limited and general character that both facilitate and require interpretation. Newer prompt-based methods that both interpret text and produce human-readable outputs still require serious work before we understand what we want to ask of them and how we want to ask for it. Additionally, smaller LMs that are computationally accessible still struggle to annotate highly subtle phenomena with extremely high precision. However, combining the expertise of subject-matter experts and traditional qualitative research methods with contemporary LMs allows for us to begin to study these complex topics on previously unreachable scales, opening possibilities for many new avenues of research.
@@ Acknowledgements
We would like to thank Erin Boell, Carolyn Chui, Estefania DeJesus, Alexandra Gier, Georgia Lawrence, Sushmi Majumder, and Elina Natarajan for their substantial annotation contributions. In addition, we thank Axel Bax, Federica Bologna, Anna Choi, Sil Hamilton, Kiara Liu, Landon Schnabel, Rosamond Thalken, Andrea Wang, Matthew Wilkens, and Shengqi Zhu for their thoughtful feedback. This work was supported in part by the NEH project AI for Humanists and the Society for the Scientific Study of Religion (SSSR) 2025 Student Research Grant.